draw the basic graph and draw a new graph for each shift
17. f(x) = 9+, (х+3)2 екуьное куска арміру
Answer #1

17)

\small f(x)=-9+\frac{1}{(x+3)^{2}}

The parent function is

\small g(x)=\frac{1}{x^{2}}\rightarrow (\textup{shift graph left by 3 units})\rightarrow g(x)=\frac{1}{(x+3)^{2}}

\small g(x)=\frac{1}{(x+3)^{2}}\rightarrow (\textup{shift graph down by 9 units})\rightarrow f(x)=-9+\frac{1}{(x+3)^{2}}

1596243580124_image.png

Therefore, graph of

\small f(x)=-9+\frac{1}{(x+3)^{2}}

1596243746869_image.png